# Synthesizing data products, mathematical models, and observational measurements for lake temperature forecasting

### v1

@@ 5 Out of sample results
Here we lay out evaluation metrics and comparators and then describe an out-of-sample exercise designed to explore the relative merits of GPBC forecasts, concluding with an analysis of results.
@@ 5.1 Validation and benchmarking
When new sensor measurements come in at midnight, for “yesterday”, we have the opportunity to assess the accuracy of our forecasts for that day, as predicted on all thirty previous days at varying future horizons. In our assessments, we record (1) RMSE as a measure of accuracy; (2) a proper log score (Eq. (25) of Gneiting and Raftery, 2007) combining accuracy and UQ out of sample; and (3) confidence interval width and (4) coverage. Given predictions $\hat{\mu}\equiv\hat{\mu}(\mathcal{X})=(\hat{\mu}_{\_}1,\dots,\hat{\mu}_{\_}N)^{\top}$ for out of sample $Y\equiv Y(\mathcal{X})=(y_{\_}1,\dots,y_{\_}N)^{\top}$ score and RMSE may be calculated as follows:
Smaller RMSEs and larger scores are preferred. When validating for multiple horizons, and/or over multiple days, we aggregate accordingly. Score, RMSE, interval width and coverage are interesting because they provide a general sense of how the forecaster is performing. However, they are most useful when benchmarking against other, competing frameworks.
As comparators we consider the following four methods. Each is simplification of flagship GPBC (with $\phi$) approach, missing one or more of its building blocks.
@@
    a GP surrogate for GLM without bias-correction (GPGLM), but otherwise using all predictors (including $\phi$) as well as Vecchia/SK features for heteroskedastic modeling and fast computation. The setup here follows Figure 10 for iterative updating, but skipping the bottom left bias-correction corner of the circuit.
  • GPO:
    a heteroscedastic GP fitted to observed sensor measurements only, using Julian day of year and depth as inputs (no $\phi$). For this competitor, we do not follow Figure 10. The resulting forecasts simply reflect a “typical” year based all previous years’ measurements. Importantly, GLM’s short-term forecasts are not used at all. In this way, GPO represents a “climatological” model, as its predictions reflect typical seasonality and general trends. Here, since the sensor data set is relatively small ($365\times 10$/year) we do not need the Vecchia approximation, so we opted for a heteroskedastic GP fit from the hetGP R package (Binois and Gramacy, 2021) via mleHetGP. A hetGP fit is similar to SK [Section 3.2], but with joint inference for first and second-moment GPs, which is crucial when the degree of replication is low. Recall that SK requires a moderate degree of replication, and there is no replication in the sensor measurements.
@@
    raw 1–30 day-ahead NOAA-GLM forecasts in the testing time period. For this competitor, we calculated empirical means and standard deviations over all 31 ensemble members for all combinations reference date, depth, and horizon in the test time period.
@@
    a bias-corrected surrogate for GLM where the input $\phi$ (providing annual variability) is omitted. Like GPO, this competitor is climatological, as its forecasts are based only on historical information. However, it does incorporate NOAA-GLM hindcasts and as forecasts for the current day, so it is not a purely historical model.
Each benchmarking comparator offers a glimpse into the potential benefit, or hazard, of higher fidelity modeling via the components that make up GPBC. GPGLM illuminates the value of bias correction, whereas GLM (which does not include a bias-correction) illustrates the value of surrogate modeling, as opposed to using raw simulations. Climatological models, like GPO and GPBC without $\phi$, are often hard to beat on longer time scales (Thomas et al., 2020), i.e., in situations where dynamics are more climate- than weather-driven. Weather forecasts become rapidly more inaccurate more than a week into the future. Although we would hope that our principal model, GPBC, would be capable of detecting and correcting for those inaccuracies, there is always estimation risk inherent in additional complexity. In other words, it could be better to simply ignore weather predictions for longer horizons.
@@ 5.2 Forecasting exercise
Our exercise simulates one year of daily forecasting (e.g. Figure 10), beginning June 11 2022, and ending on the same day in 2023, treating a corpus of model runs and sensor measurements from October 3 2020 through June 10 2022 as training data which is augmented daily. GPBC and GPGLM model fitting is as described in Section 3. Training data for GPO consisted only of sensor measurements comprising the inital training period before June 10 2022. We present our results in several views.
@@
Figure 12 shows RMSE, score, coverage, and PI width for each horizon, aggregating over all depths and days in the forecasting period. First consider RMSE in panel (A), where lower is better. Observe that every method outperforms OGP, the purely sensor-based climatological benchmark, at short horizons. This is a testament to the value of NOAA-GLM in the short term. However, OGP outperforms every other method after about two weeks out. We suspect this is because NOAA forecasts are unreliable at long horizons. In the shorter term, those methods that are most influenced by GLM forecasts perform best. Standalone GLM forecasts or GPGLM (with the surrogate) have similar performance. Improvements in accuracy due to the bias-correction in GPBC (with $\phi$) is clearly evident, being consistently better than both GPGLM and GLM. However, at more than two weeks out, climatological forecasts (OGP) are most accurate. Omitting $\phi$ compromises the influence of GLM forecasts, which results in some advantages as well as disadvantages. Over short horizons, this comparator underwhelms, but at longer horizons its forecast are more robust, similar to OGP.
@@
Next consider the panel (B) of the figure, which summarizes scores over horizon (higher is better). Overall trends here are similar to the RMSE panel, with some notable exceptions. The raw GLM’s score falls outside the lower range of the $y$-axis, so a line for this comparator does not appear on the plot. Although it was reasonably accurate in terms of RMSE (left panel), it has poor UQ, which we shall discuss next with panels (D–E). Again, our flagship GBPC comparator performs best for the first two weeks, except horizon one, where the surrogate-assisted GLM (GPGLM) has a better score than GPBC. However, like raw GLM, this comparator is exceedingly poor at higher horizons. Observe that GPBC without $\phi$ is the best comparator for horizons longer than two weeks. In particular, it beats the purely sensor-based climatological comparator.
@@
Panel (C) summarizes empirical coverage from 90% predictive intervals over horizon. Only the methods using surrogate-assisted NOAA-GLM forecasts can achieve nominal coverage (at 90%) at any horizon. Raw-GLM dramatically under-covers. It’s surrogate (GPGLM), by contrast, is able to correctly quantify uncertainty by accessing the entire corpus of simulations across the years. Although empirical coverage drops off from nominal at higher horizons, models that incorporate GLM forecasts (GPBC and GPGLM) always outperform OGP, the climatological benchmark. Additional insight into these coverage results is provided by panel (D). Observe that the PIs of raw GLM forecasts are much narrower than those from other competitors, especially at early horizons. We also know (e.g., Figure 8) that standalone GLM simulations and forecasts can exhibit pronounced bias. Taken together, this explains raw GLM’s poor scores (panel B). Also note that PIs resulting from GPGLM are twice as wide as those of GLM for all horizons. Although GPBC and GPGLM have similar PI coverage, PIs of GPBC are much narrower than those of GPGLM. In fact, at horizon 30, PI width of GPBC (2.2) is nearly half that of GPGLM (4.2). Together, results show that bias-correction not only increases accuracy (via RMSE in panel A) but also results in narrower PIs. Note that although GPBC utilizes a homoskedastic GLM surrogate, PI interval width expands only slightly over horizon because the bias correction is homoskedastic.
@@
Figure 14: Forecasts from GPGLM (blue) and GPBC (orange) generated during the forecasting exercise for horizon 1 (one day into the future, top) and horizon (30, 30 days into the future) over day of year (DOY) at surface, 2, 4, 6, and 8m depths (each panel denotes one depth). Observations are shown in purple.
@@
Figure 14 offers another perspective by focusing on forecasts over day of year (DOY) for two forecast horizons (1 and 30 days into the future). The top panel shows one-day ahead forecasts generated by GPGLM and GPBC, while the bottom panel shows forecasts at 30 days into the future at five depths. Observe in the top panels (one-day ahead) that GPGLM and GLM are generally in agreement with observations. In fact, for the majority of days of the year and all depths, forecasts and observations lie nearly on top of one another. At thirty days out (bottom panels), the discrepancies between forecasts from GPGLM and GPBC are much more apparent, especially for deeper depths. At surface depths, the effect of bias-correction is not as obvious, but the improvement from bias-correction clearly results in enhanced forecast accuracy at deeper depths between DOY 1 and 240. Taken together with main results over depth and horizon in Figures 1213, the visual here reinforces the importance of a bias-correction in improving accuracy over depth and horizon.
@@
Figure 15: Bias (e.g., $\bar{\mathbf{Y}}_{\_}n^{F}-\hat{\mu}_{\_}n^{\mathrm{SK}}(\bar{\mathbf{X}}_{\_}n)$ for out-of sample forecasts from the GLM surrogate (pink), raw GLM (gray) and predicted bias (teal), averaged over depth and horizon for dates in the testing time period. Vertical black dashed lines denote example DOYs depicted in Figure 16.
@@
Next, Figure 15 shows how raw GLM bias (gray) surrogate GLM bias (“actual bias” via $\bar{\mathbf{Y}}_{\_}n^{F}-\hat{\mu}_{\_}n^{\mathrm{SK}}(\bar{\mathbf{X}}_{\_}n)$), and predicted bias vary over DOY in our out-of-sample forecasting exercise. Observe that while accuracy varies over horizon [Figure 12] the magnitude of bias from GLM depends even more strongly on the time of year a forecast is made [Figure 15]. Bias from GLM and its surrogate are similar. Predicted bias, which is used to correct it, follows a similar pattern, but it is not perfectly aligned with actual bias: it underestimates for the first half of the year, and then overestimates in the second half. Note that the predicted bias is not smooth, because bias for each day in the forecasting period was predicted iteratively (following Figure 10), so each prediction was based upon a slightly different GP fit.
Finally, Figure 16 augments the visual in Figure 15 by breaking out horizon and depth contributions to the bias for three particular days (indicated as vertical dashed lines in Figure 15). These three example days were chosen to display occasions when the bias-correction was correct in sign but not in magnitude (DOY 48 and 260) and generally correct in sign and magnitude (171). Consider DOY 48 first, in the left panels of Figure 16. Applying a bias correction (GPBC) improves the accuracy of forecasts compared to the GLM surrogate without a bias correction (GPGLM) with respect to observations (purple dots) at both 0 and 5m depths, although at 5m depth the forecast is still characterized by a cold bias. Moving to DOY 260 (right panel), the bias-correction actually worsens the forecast at the surface depth at after 15 days into the future. Finally, at day 171 (middle panel), forecasts at surface and 5m depths are generally improved by the bias-correction. Though we only have about 2.5 total years of training data, these plots demonstrate that predicted bias applied to out-of-sample forecasts from GPBC mainly improves forecasts, although there are periods during the year when bias-correction over- or under-corrects or is of the wrong magnitude.
6 Discussion
@@
We introduced an innovative method for forecasting lake temperature profiles 1 to 30 days into the future, bias correcting and surrogate modeling a computer simulation campaign using a lake ecosystem model, GLM. Our framework is thrifty in spite of massive data and cubic bottlenecks, enables iterative forecasting, accounts for input-dependent variance. When compared to other methods, without all of these ingredients, forecasts produced by our bias-corrected Gaussian process surrogate (GPBC) are typically more accurate and deliver superior uncertainty quantification (UQ), especially for forecasts up to two weeks ahead. However, climatological forecasts remain more accurate for predictions beyond two weeks, and in our future work we intend to explore possible reasons for this.
We suspect the main reason we lose to the climatological model (GPO) is a paucity of training data for estimating bias and the quality of NOAA forecasts used to drive GLM. We only have 2.5 years, meaning that we’ve seen fewer than three examples of yearly variation. Although GPO has access to the same amount of observational data, its simpler design – not requiring the correction of poor model-based forecasts at longer horizons – helps to prevent over-fitting. Moreover, because the variability of GLM ensemble forecasts increases with forecast horizon, it is likely that the shape and degree of bias at longer horizons is more difficult to estimate compared to shorter horizons, where ensemble forecasts are more in agreement. Thus, bias correction at longer horizons may be less effective than at lower horizons without many more examples of yearly variation.
The quality of NOAA-GLM forecasts is largely dependent on the quality of NOAA forecasts themselves (assuming that GLM is well-calibrated). Weather forecast accuracy more than 10 days into the future declines quickly (Bauer, Thorpe and Brunet, 2015). Moreover, NOAA forecasted variables represent spatial averages over a 25km grid cell, much bigger than FCR for example, which may be too coarse to capture fine-scale variation in meteorological variables. Downscaling and bias-correction of NOAA forecasted variables could result in more accurate lake temperature forecasts from GLM, but additional post-processing of NOAA forecasts would take substantial computation time. Consequently, we decided that and investigation was beyond the scope of our current study. While forecasts of the climatological benchmark model were more accurate than GPBC at longer horizons, the UQ of those forecasts were generally poorer compared to those of GPBC.
Another way to extend this work would be to use GPBC as a likelihood for setting configuration parameters to GLM. In other words, to use the surrogate as a statistical calibration mechanism for the computer model (Kennedy and O’Hagan, 2001; Liu et al., 2009), as opposed to relying on expert opinion using observed meteorological data, not NOAA forecasts. We could also treat the derived input $\phi(t,\lambda)$, allowing us to distinguish forecasts among years, (i.e., the weather inputs) as a latent variable that could be estimated along with other quantities, thus allowing us to distinguish forecasts among years. Such an approach, which involves estimating a high-dimensional free quantity for all $t$ and $\lambda$, is precluded, at the moment, by a paucity of sensor measurements. Perhaps once we have accumulated 10+ years of additional data , spanning more yearly environmental variability, the benefits of such an approach might outweigh the additional estimation risk it implies.
@@
Appendix A Toy problem
This 1d example was introduced by Binois, Gramacy and Ludkovski (2018). The mean function is $f(x)=2\times\mathrm{exp}(-30(x-0.25)^{2}+\mathrm{sin}(\pi x^{2}))-2$ with a noise function of $r(x)=\frac{1}{3}\mathrm{exp}(\mathrm{sin}(2\pi x))$. Observations are simulated as $y\sim f(x)+\epsilon$, where $\epsilon\sim N(0,\sigma^{2}=r(x)).$ The training data comprise of 100 input locations equally spaced in $[0,1]$, with 15 replicates observed upon each. Therefore, the total number of observations is $N=15\times 100=1500$. The simulated data can be represented as $D_{\_}N=(x_{\_}i,y_{\_}i),i=1\dots N$. Let $\mathbf{X}_{\_}N$ and $\mathbf{Y_{\_}N}$ collect inputs and outputs, respectively. We then model $D_{\_}N$ as $\mathcal{GP}(D_{\_}N)$ and use four fitted modes:
  1. 1.

### v2

@@ 5 Out of sample results
Here we lay out evaluation metrics and comparators and then describe an out-of-sample exercise designed to explore the relative merits of GPBC forecasts.
@@ 5.1 Validation and benchmarking
When new sensor measurements come in at midnight we have the opportunity to assess the accuracy of yesterday’s forecasts, as predicted on all thirty previous days at varying future horizons. We evaluate (1) RMSE; (2) a proper log score (Eq. (25) of Gneiting and Raftery, 2007); (3) confidence interval width and (4) coverage. For details see Appendix C. Smaller RMSEs and larger scores are preferred.
As comparators we consider the following four methods. Each is simplification of our flagship GPBC (with $\phi$) approach, missing one or more of its building blocks.
@@
    a GP surrogate for GLM without bias-correction (GPGLM), but otherwise using all predictors (including $\phi$) and Vecchia/SK for fast heteroskedastic modeling.
  • OGP:
    a heteroscedastic GP fitted to observed sensor measurements only, using Julian day of year and depth as inputs (no $\phi$). These forecasts simply reflect a “typical” year based all previous years’ measurements. Importantly, GLM’s short-term forecasts are not used at all. In this way, OGP represents a “climatological” model, as its predictions reflect typical seasonality and general trends. Since the sensor data set is smaller ($365\times 10$/year) we do not need Vecchia, so we opted for a heteroskedastic GP fit from hetGP (Binois and Gramacy, 2021) via mleHetGP for joint inference for first and second-moment GPs.
@@
    raw 1–30 day-ahead NOAA-GLM forecasts in the testing time period. We calculated empirical means and standard deviations over all 31 ensemble members for all combinations reference date, depth, and horizon in the test time period.
@@
    a bias-corrected surrogate for GLM where the input $\phi$ omitted. Like OGP, forecasts are based only on historical information. However, it does incorporate NOAA-GLM hindcasts for the current day, so it is not purely climatological.
Each benchmarking comparator offers a glimpse into the potential benefit, or hazard, of higher fidelity modeling via the components that make up GPBC. GPGLM illuminates the value of bias correction, whereas GLM (which does not include a bias-correction) illustrates the value of surrogate modeling, as opposed to using raw simulations. Climatological models, like OGP and GPBC without $\phi$, are often hard to beat on longer time scales (Thomas et al., 2020), i.e., in situations where dynamics are more climate- than weather-driven. Weather forecasts become rapidly more inaccurate more than a week into the future.
@@ 5.2 Forecasting exercise
We simulated one year of daily forecasting beginning June 11 2022 treating a corpus of model runs and sensor measurements from October 3 2020 through June 10 2022 as training, augmented daily. GPBC and GPGLM model fitting is as described in Section 3. Training data for OGP consisted only of sensor measurements comprising the inital training period before June 10 2022. We present our results in several views.
@@
Figure 9 shows RMSE, score, coverage, and PI width for each horizon, aggregating over all depths and days in the forecasting period. First consider RMSE in panel (A), where lower is better. Observe that every method outperforms OGP, the purely sensor-based climatological benchmark, at short horizons. This is a testament to the value of NOAA-GLM in the short term. However, OGP outperforms every other method after about two weeks out. We suspect this is because NOAA forecasts are unreliable at long horizons. In the shorter term, those methods that are most influenced by GLM forecasts perform best. Standalone GLM forecasts or GPGLM (with the surrogate) have similar performance. Improvements in accuracy due to the bias-correction in GPBC (with $\phi$) is clearly evident, being consistently better than both GPGLM and GLM. However, after two weeks climatological forecasts (OGP) win. Omitting $\phi$ compromises the influence of GLM forecasts, which results in some advantages as well as disadvantages. Over short horizons, this comparator underwhelms, but at longer horizons its forecasts are more robust, similar to the purely climatological model OGP.
@@
Next, consider panel (B), which summarizes scores over the horizon (higher is better). Overall trends here are similar to the RMSE panel, with some notable exceptions. The raw GLM’s score falls outside the lower range of the $y$-axis, so a line for this comparator does not appear on the plot. Although it was reasonably accurate in terms of RMSE (left panel), it has poor UQ, which we shall discuss next with panels (D–E). Again, our flagship GBPC comparator performs best for the first two weeks, except horizon one, where the surrogate-assisted GLM (GPGLM) has a better score than GPBC. However, like raw GLM, this comparator is poor at higher horizons. Observe that GPBC without wins for horizons longer than two weeks. In particular, it beats the purely sensor-based climatological competitor OGP.
@@
Panel (C) summarizes empirical coverage from 90% predictive intervals over horizon. Only the methods using surrogate-assisted NOAA-GLM forecasts can achieve nominal coverage (at 90%) at any horizon. Raw-GLM dramatically under-covers. Its surrogate (GPGLM), by contrast, is able to correctly quantify uncertainty by accessing the entire corpus of simulations across the years. Although empirical coverage drops off from nominal at higher horizons, models that incorporate GLM forecasts (GPBC and GPGLM) always outperform OGP, the climatological benchmark. Additional insight into these coverage results is provided by panel (D). Observe that the PIs of raw GLM forecasts are much narrower than those from other competitors, especially at early horizons. We also know (e.g., Figure 6) that standalone GLM simulations and forecasts can exhibit pronounced bias. Taken together, this explains raw GLM’s poor scores (panel B). Also note that PIs resulting from GPGLM are twice as wide as those of GLM for all horizons. Although GPBC and GPGLM have similar PI coverage, PIs of GPBC are much narrower than those of GPGLM. In fact, at horizon 30, PI width of GPBC (2.2) is nearly half that of GPGLM (4.2).
@@
Figure 10: Forecasts from GPGLM (blue) and GPBC (orange) generated during the forecasting exercise for horizon 1 (one day into the future, top) and horizon (30, 30 days into the future, bottom) over day of year (DOY) at surface, 2, 4, 6, and 8m depths (each panel denotes one depth). Observations are shown in darker purple dots.
@@
Appendix D.1 provides a similar analysis over depth. Figure 10 focuses on forecasts over day of year (DOY) for two forecast horizons (1 and 30 days into the future). The top panel shows one-day ahead forecasts generated by GPGLM and GPBC, while the bottom panel shows forecasts at 30 days into the future at five depths. Observe in the top panels (one-day ahead) that GPGLM and GLM are generally in agreement with observations, lying nearly on top of one another. At thirty days out (bottom panels), the discrepancies between forecasts from GPGLM and GPBC are much more apparent, especially for deeper depths. At surface depths, the effect of bias-correction is not as obvious, but the improvement from bias-correction enhances accuracy at deeper depths between DOY 1 and 240.
@@
Figure 11: Bias (e.g., $\bar{\mathbf{Y}}_{\_}n^{F}-\hat{\mu}_{\_}n^{\mathrm{SK}}(\bar{\mathbf{X}}_{\_}n)$ for out-of sample forecasts from the GLM surrogate (solid pink line), raw GLM (solid gray line) and predicted bias (dot-dash teal line), averaged over depth and horizon for dates in the testing time period. Vertical black dashed lines denote example DOYs depicted in Figure 16.
@@
Next, Figure 11 shows how raw GLM bias (gray) surrogate GLM bias (“actual bias” via $\bar{\mathbf{Y}}_{\_}n^{F}-\hat{\mu}_{\_}n^{\mathrm{SK}}(\bar{\mathbf{X}}_{\_}n)$), and predicted bias vary over DOY in our out-of-sample forecasting exercise. Observe that while accuracy varies over horizon [Figure 9] the magnitude of bias from GLM depends even more strongly on the time of year a forecast is made [Figure 11]. Bias from GLM and its surrogate are similar. Predicted bias, which is used to correct it, follows a similar pattern, but it is not perfectly aligned with actual bias: it underestimates for the first half of the year, and then overestimates in the second half. Note that the predicted bias is not smooth, because bias for each day in the forecasting period was predicted iteratively, so each prediction was based upon a slightly different GP fit. Appendix D.2 provides an additional set of visuals breaking out horizon and depth contributions to bias for three particular days.
6 Conclusion
@@
We studied forecasting lake temperatures 1 to 30 days into the future, bias correcting and surrogate modeling a computer simulation campaign using a lake ecosystem model, GLM. Our framework is thrifty in spite of massive data and cubic bottlenecks, enables iterative forecasting, accounts for input-dependent variance. When compared to other methods, without all of these ingredients, forecasts produced by our bias-corrected Gaussian process surrogate (GPBC) are typically more accurate and deliver superior uncertainty quantification (UQ), especially for forecasts up to two weeks ahead. However, climatological forecasts remain more accurate for predictions beyond two weeks, and in our future work we intend to explore possible reasons for this. Further discussion and additional thoughts for future work are provided in Appendix E.
7 Data and code availability
Code to replicate all modeling and figures is available in a Github repo: https://github.com/maikeh7/Surrogate_Assisted_GLM. All data may be found in a Zenodo repo: https://zenodo.org/uploads/10028017.
@@ 7 Data and code availability
{supplement}
Appendix A Vecchia approximation
Here we provide detail and empirical work supporting the discussion in Section 3.3.
A.1 Review
The crux of the Vecchia approximation is a simple identity linking joint and conditional distributions. Consider a vector $\mathbf{Y}_{\_}n$ of responses, like one filled with $\bar{y}_{\_}i$-values in Section 3.2. The joint probability of $\mathbf{Y}_{\_}n$ can be factorized as
Some additional implementation details make things even faster and allow for distributed computation, which we leave to our references (Stein, Chi and Welty, 2004; Guinness, 2018; Katzfuss and Guinness, 2021; Stroud, Stein and Lysen, 2017; Datta et al., 2016; Wu, Pleiss and Cunningham, 2022; Katzfuss et al., 2020a). We have been satisfied with the performance of the defaults offered by the R referenced in Section 3.3, synthesizing many of those elements including composition of the $c(i)$. We settled on $m=30$ by entertaining an out-of-sample prediction exercise similar to that reported by Sauer, Cooper and Gramacy (2023), which showed diminished returns for larger $m$-values. Perhaps the most important aspect of that software, compared to other methods based on the Vecchia approximation, is that the effect of an input-dependent lengthscale $\boldsymbol{\gamma}$ can be learned along with scale $\tau^{2}$ and (scalar) nugget $g$ via Fisher scoring (Guinness, 2021). The authors call this the “Scaled Vecchia” approximation; however we’ll simply refer to it as “Vecchia” with the understanding that input-dependent lengthscales are being learned. With this setup we are able to fit GPs with $n\approx 270{,}000$ as described in Section 3.2 in about twenty seconds on an ordinary workstation (64-bit Intel(R) Core(TM) i9-9900K CPU @ 3.60GHz with 16 cores).
A.2 Toy problem
This 1d example was introduced by Binois, Gramacy and Ludkovski (2018). The mean function is $f(x)=2\times\mathrm{exp}(-30(x-0.25)^{2}+\mathrm{sin}(\pi x^{2}))-2$ with a noise function of $r(x)=\frac{1}{3}\mathrm{exp}(\mathrm{sin}(2\pi x))$. Observations are simulated as $y\sim f(x)+\epsilon$, where $\epsilon\sim N(0,\sigma^{2}=r(x)).$ The training data comprise of 100 input locations equally spaced in $[0,1]$, with 15 replicates observed upon each. Therefore, the total number of observations is $N=15\times 100=1500$. The simulated data can be represented as $D_{\_}N=(x_{\_}i,y_{\_}i),i=1\dots N$. Let $\mathbf{X}_{\_}N$ and $\mathbf{Y_{\_}N}$ collect inputs and outputs, respectively. We then model $D_{\_}N$ as $\mathcal{GP}(D_{\_}N)$ and use four fitted models:
  1. 1)